# ShortcutsBench: A Large-Scale Real-world Benchmark for API-based Agents

### v2

@@ Abstract
Recent advancements in integrating large language models (LLMs) with application programming interfaces (APIs) have gained significant interest in both academia and industry. These API-based agents, leveraging the strong autonomy and planning capabilities of LLMs, can efficiently solve problems requiring multi-step actions. However, their ability to handle multi-dimensional difficulty levels, diverse task types, and real-world demands through APIs remains unknown. In this paper, we introduce ShortcutsBench, a large-scale benchmark for the comprehensive evaluation of API-based agents in solving tasks with varying levels of difficulty, diverse task types, and real-world demands. ShortcutsBench includes a wealth of real APIs from Apple Inc.’s operating systems, refined user queries from shortcuts, human-annotated high-quality action sequences from shortcut developers, and accurate parameter filling values about primitive parameter types, enum parameter types, outputs from previous actions, and parameters that need to request necessary information from the system or user. Our extensive evaluation of agents built with $5$ leading open-source (size >= 57B) and $4$ closed-source LLMs (e.g. Gemini-1.5-Pro and GPT-3.5) reveals significant limitations in handling complex queries related to API selection, parameter filling, and requesting necessary information from systems and users. These findings highlight the challenges that API-based agents face in effectively fulfilling real and complex user queries. All datasets, code, and experimental results will be available at https://github.com/eachsheep/shortcutsbench.
@@ 1 Introduction
Large language model based agents (LLM-based agents) (Wang et al., 2023b; Xi et al., 2023) built on application programming interfaces (APIs) (Qin et al., 2023; Huang et al., 2023) have recently gained significant interest in both academia (Shen et al., 2024; Wang et al., 2023b) and industry (Microsoft, 2024; OpenAI, 2024c). By integrating LLM with APIs, these agents can access real-time information (OpenAI, 2024a; Microsoft, 2024), reduce hallucination with external knowledge (Li et al., 2023a; Gao et al., 2023), and automatically plan and complete complex tasks that need multi-step actions (Gravitas, 2024; Pan et al., 2023). Many of these agents (OpenAI, 2024c; Microsoft, 2024; Gravitas, 2024) have also demonstrated commendable performance on simple tasks involving only a few actions such as “Check the weather ① and tell me ②” (OpenAI, 2024c). These impressive performances raise an important question: Do these API-based agents truly possess the capability to generate complex action sequences for real demands with real APIs?
@@
Some existing benchmarks / datasets (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024) have attempted to evaluate API-based agents. However, they have three limitations: First, the APIs (a.k. tools available to the agent) lack richness, and the queries (a.k. the task to the agent) lack complexity (Table 1). They either involve a limited number of APIs (Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024) or cover only a few numbers of apps (Patil et al., 2023; Tang et al., 2023; Schick et al., 2024; Hao et al., 2024) (an app may have one or multiple APIs), or the difficulties of the queries are limited in a narrow range, with the average action length ranges from $1$ to $5.9$ (Table  1). This lack of richness and complexity makes it difficult to effectively distinguish the capabilities of different agents, particularly on larger and smarter LLMs like Gemini-1.5 Pro DeepMind (2024) and GPT-3.5-turbo OpenAI (2023a). Second, the APIs lack realism as they may be manually crafted, and the queries fail to reflect actual user demands since they may be either created by hand or generated directly by ChatGPT OpenAI (2023a) without verifying real user demands (Table 1). Moreover, they only cover the evaluation of API selection, lacking a study on API parameter filling (Table 1). Efficient and accurate parameter filling is essential for an agent to complete tasks successfully. Thirdly, they don’t adequately evaluate the agent’s ability to ask systems or the users for the necessary information to resolve the missing information for solving the queries (Table 1). This is crucial as a user’s query may be implicit or may not provide all the information an agent needs to solve the task (Qin et al., 2023; Qian et al., 2024) effectively.
@@
In this paper, we present ShortcutsBench. To our best known, ShortcutsBench is the first large-scale real API-based agent benchmark considering APIs, queries, and corresponding action sequences. ShortcutsBench provides rich real APIs, queries with various difficulties and task types, high-quality human-annotated action sequences by shortcuts developers, along queries from real user demands. Moreover, it also provides precise values for parameter filling, including primitive data types, enum types, and the use of output from previous actions for parameter values, as well as evaluations of the agent’s awareness in requesting necessary information from the system or user. Furthermore, the scale of APIs, queries, and the corresponding action sequences in ShortcutsBench is comparable or even better to benchmarks and datasets created by LLM or modified by existing datasets. The overall comparison between ShortcutsBench and existing benchmarks / datasets is listed in Table 1.
We conducted extensive evaluations of API-based agents from $9$ leading LLMs on  ShortcutsBench, including the evaluation of API selection, parameter value filling, and recognition of the need for input from the system or the user. The chosen LLMs including four closed-sourced LLMs like Gemini-1.5-Pro (DeepMind, 2024) and GPT-3.5-turbo (OpenAI, 2024b), and five open-source LLMs like LLaMA-3-70B (Meta, 2024) and QWen-2-70B (Qwen, 2024b). Our findings highlight the limitations of these agents in addressing real, rich, and complex user queries. In summary, this paper makes the following key contributions:
  • To our best known, we have built the most realistic, rich, and comprehensive API-based agent benchmark. This benchmark is even comparable in scale to existing benchmarks / datasets built using LLMs (Table 1).
@@
  • We evaluated nine most advanced and mainstream LLM-based agents on all operations required to complete user queries, including API selection, parameter filling, and their awareness to request necessary information from the system or user when needed.
@@
  • We obtained massive interesting conclusions such as (1) The performance gap between open-source and closed-source LLMs has become very small; (2) Existing LLMs still have significant shortcomings in multi-step reasoning; (3) Extracting necessary parameters from queries is the most challenging task in parameter filling; (4) There is a substantial lack of awareness in agents when it comes to requesting necessary information.
@@
Table 1: ShortcutsBench has a great advantage in the realness, richness, and complexity of APIs, queries, and corresponding action sequences, the validity of action sequences, accurate parameter value filling, the awareness for asking information from the system or the users, and the overall scale.
@@
# Apps 88 390 3451 3 53 400 8 13
# APIs 1414 390 16464 1645 53 400 232 13
# Queries 7627 21127 12657 17002 3938 274 2726 1530
# Avg APIs 9.62 1.0 2.3 1.0 1.0 2.1 5.4 13.0
# Avg Actions 21.62 1.0 4.0 1.0 1.0 2.2 5.9 1.0
@@ 2 Related Work
API-based agents treat APIs as tools (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024; Zhu et al., 2023; Gravitas, 2024; AgentGPT, 2023). They accept queries, generate action sequences based on queries and provided APIs, and generate next action depends on the history actions (Yao et al., 2022). Related work about API-based agents can generally be categorized into $3$ types depending on the objective: (1) Task-specific enhancement focuses on improving the agent’s ability to solve a specific type of task like game and question-answering (Hao et al., 2024; Zhu et al., 2023; Gravitas, 2024; AgentGPT, 2023). (2) Data-driven workflows emphasize the importance of data by researching how to construct workflows to get action sequences, enabling generated data to fine-tune the model (Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024). (3) Agent evaluation studies the assessment of agents (Huang et al., 2023; Li et al., 2023b).
@@
Code-based agents use code generated as a medium for interaction with the external environment (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023). They accept queries, generate scripts in programming languages such as Python (OpenAI, 2023b), JavaScript (Wang et al., 2023a), or Shell (OpenInterpreter, 2024), and then input the code into interpreters. The results of the interpreter are then returned to the agent, which is used to help determine the next action in code generation. Currently, these approaches primarily focus on enhancing agent performance in specific tasks by incorporating additional knowledge (Wang et al., 2023a; Wu et al., 2023), increasing feedback (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023), and decomposing tasks (OpenAI, 2023b; Wang et al., 2023a; OpenInterpreter, 2024; Wu et al., 2023).
@@
DAPs (Abdou et al., 2021; Coze, 2023; Rahmati et al., 2017; Chakraborti et al., 2020) refer to software tools or services designed to optimize workflows through automation. DAPs leverage technologies such as robotic process automation (RPA) (Chakraborti et al., 2020) and low-code/no-code development tools to achieve the goals. DAPs like Zapier (Abdou et al., 2021; Rahmati et al., 2017), Make (Make, 2023), and IFTTT (Abdou et al., 2021; Rahmati et al., 2017) offer extensive APIs that enable users to create automated workflows. Similarly, DAPs such as Microsoft Power Automate (Abdou et al., 2021) and Tasker (Dias, 2024) are primarily used to build workflows on Azure and Android, respectively. Recently, with the rise of LLM-based agents, platforms like Coze (Coze, 2023) and Dify (Dify, 2023) have emerged as “agent construction platforms”. Functionality like “workflow” in these platforms can also help manually build workflows, but they have been specifically optimized for integration with LLMs.
Shortcuts app (formerly Workflow(Apple, 2024a) is an app developed by Apple for building workflows through a graphical interface, available on Apple’s operating systems (iOS/iPadOS and macOS). Shortcuts app can be seen as the DAP of Apple. It allows users to create workflows (known as shortcuts (Apple., 2024)) that execute specific tasks on their devices and share them online via iCloud (Apple, 2024c). Users can also download curated shortcuts from the Gallery of the Shortcuts app. However, the shortcuts available in the Gallery are very limited, with only a few dozen options. To access more shortcuts, users must either collect them from third-party sharing sites like Shortcuts Gallery (Gallery, 2024) and SSPAI (SSPai, 2024) or create their own. Shortcuts can be triggered through the Shortcuts app, widgets, the share sheet, and Siri, and they can also be automated to run upon specific events.
Shortcuts are composed of multiple API calls (actions). An agent can use the shortcut as a whole API or utilize the individual APIs involved in the shortcut. This paper treats the various APIs within the shortcuts as APIs available to the agent, aiming for the agent to automatically construct workflows of API calls.
@@ 3 Dataset
In this Section, we first introduce the acquisition of the dataset (Section 3.1). Then, we outline the ShortcutsBench’s construction process (Section 3.2). Finally, we outline the setup for evaluation tasks to evaluate the agent’s ability to handle tasks of varying difficulty (Section 3.3.1), including the ability to select suitable APIs, the ability to do parameter filling (Section 3.3.2), and the awareness in requesting additional information from the system or user (Section 3.3.3).
@@ 3.1 Dataset Acquisition
Figure 1: (a) illustrates the data acquisition process. (b) shows the dataset acquisition of existing work (Table 1). APIs in existing benchmarks / datasets are created by hand or modified from existing datasets, or API hubs, and queries and action sequences are constructed using templates or semi/fully automated methods with LLMs.
Figure 1 shows the data acquisition process. We first use search engines to identify popular public shortcut-sharing sites ①. We totally find $14$ sites such as Shortcuts Gallery (Gallery, 2024) and SSPai (SSPai, 2024). Then we crawled these sites to obtain fields such as “shortcut name”, “function description”, “shortcut type”, and “iCloud link” ②. After deduplicating based on “iCloud link” (Apple, 2024c), we got the source files of all $8675$ shortcuts (Apple, 2024b) ③. Subsequently, we extracted “app name” using the  field WFWorkflowActionIdentifier in the source file like com.openai.chat.AskIntent, and then downloaded related apps ④ from various sources: (1) third-party apps from the “macOS App Store” or the official website of the app, (2) non-system apps (uninstallable) like Keynote from path /Applications/ on macOS, (3) system apps like Reminders (uninstallable) from path /System/Application/ on macOS, (4) iOS apps from the “iOS App Store”, (5) Shortcuts app itself from path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. During the downloading, we also excluded some legacy apps and $12$ paid apps. For more details about the whole acquisition process, please refer to Appendix A.1.
Then we managed to extract APIs from the downloaded apps ⑤. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Apple-Inc., 2024b; 2022; 2023) framework, \${filename}.intentdefinition from SiriKit (Apple-Inc., 2024e; 2022; 2023; c) framework, and WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. We extracted all APIs involved in the app’s shortcuts. During the extraction, we perform deduplication of APIs based on manually crafted rules. Finally, as shown in Table 1, we get $88$ apps from various categories such as “Health & Fitness”(iTunes App Store, 2024b) and “Developer Tools” (iTunes App Store, 2024a). These apps include $1414$ APIs, including all of $556$ APIs involved in $7627$ shortcuts. For more details about the extraction, please refer to the Appendix A.1.

### v3

@@ Abstract
Recent advancements in integrating large language models (LLMs) with application programming interfaces (APIs) have gained significant interest in both academia and industry. Recent work demonstrates that these API-based agents exhibit relatively strong autonomy and planning capabilities. However, their ability to handle multi-dimensional difficulty levels, diverse task types, and real-world demands remains unknown. In this paper, we introduce ShortcutsBench, a large-scale benchmark for the comprehensive evaluation of API-based agents in solving real-world complex tasks. ShortcutsBench includes a wealth of real APIs from Apple Inc., refined user queries, human-annotated high-quality action sequences, detailed parameter filling values, and parameters requesting necessary input from the system or user. We revealed how existing benchmarks / datasets struggle to accommodate the advanced reasoning capabilities of existing more intelligent LLMs. Moreover, our extensive evaluation of agents built with $5$ leading open-source (size $\geq$ 57B) and $5$ closed-source LLMs (e.g. Gemini-1.5-Pro and GPT-4o-mini) with varying intelligence level reveals significant limitations of existing API-based agents in the whole process of handling complex queries related to API selection, parameter filling, and requesting necessary input from the system and the user. These findings highlight the great challenges that API-based agents face in effectively fulfilling real and complex user queries. All datasets, code, experimental logs, and results are available at https://github.com/EachSheep/ShortcutsBench .
@@ 1 Introduction
Table 1: Less intelligent LLMs (even $3B$) on existing benchmarks / dataset demonstrated excellent results with the same prompt in Section 4.1.
Acc. (%) MetaTool ToolLLM ToolBench
2024b 2024 2024
LLaMA-3.2-3B 89.64 72.92 79.47
QWen-2.5-3B 88.29 77.86 91.35
LLaMA-3-8B 89.00 78.31 93.57
QWen-2.5-7B 92.50 82.69 94.26
GPT-4o-mini 88.31 84.50 89.90
@@
Large language model based agents (LLM-based agents) (Wang et al., 2024b; Xi et al., 2023) built on application programming interfaces (APIs) have gained significant interest in academia and industry. By integrating LLM with APIs, LLMs can access real-time information (Qin et al., 2024), reduce hallucination with external knowledge (Gao et al., 2024), as well as plan and complete complex tasks that need multi-step actions (Gravitas, 2024). Many of these agents (OpenAI, 2024a) have already demonstrated commendable performance on simple tasks involving only a few actions such as “Check the weather ① and tell me ②”. This impressive performance raises an important question: Are these API-based agents truly capable of generating action sequences for real and complex demands?
@@
Some existing benchmarks / datasets [^1] We refer to the evaluation-specific datasets as ”benchmarks” and fine-tuning datasets as ”datasets”. have attempted to evaluate API-based agents. However, they have three limitations (please refer to Table 2 for all details): First, the APIs (a.k.a tools available to the agent) lack richness, and the queries (a.k.a the task to the agent) lack complexity. They either involve a limited number of APIs, cover small numbers of apps (an app may have $\geq 1$ APIs), or the difficulties of the queries are limited in a narrow range, with the average action length ranges from $1$ to $5.9$. This lack of richness and complexity makes it difficult to effectively distinguish the capabilities of different agents, even on less intelligent LLMs like QWen-2.5-3B (Alibaba, 2024), let alone more intelligent LLMs like Gemini-1.5-Pro (Google, 2024). Our evaluation on API selection of these less intelligent LLMs on $3$ representative [^2] MetaTool uses the native GPT API, while ToolBench and ToolLLM have the longest average action length and the largest scale with real-world API, respectively. benchmarks / datasets (Table  1) shows that even 3B LLMs can achieve impressive results. There is almost no difference in accuracy across LLMs of varying intelligence levels. Therefore, existing benchmarks / datasets struggle to accommodate more intelligent LLMs and to differentiate the intelligence levels among various LLMs. Second, the APIs lack realism as they may be manually crafted, and the queries fail to reflect actual user demands since they may be either created by hand or generated directly by LLMs without verifying real user demands. Moreover, they only cover the evaluation of API selection, lacking a study on API parameter filling. Efficient and accurate parameter filling is essential for an agent to finish the whole process of completing queriess Third, they don’t adequately evaluate the agent’s ability to request systems or the users for the necessary input to resolve the missing information for solving the queries. This is crucial as a user’s query may be implicit or may not provide all the input an agent needs to solve the task effectively.
@@
In this paper, we innovatively propose to use data extracted from existing Digital Automation Platforms (DAPs), Apple Shortcuts, to construct a high-quality benchmark for API-based agents, i.e., ShortcutsBench. To the best of our knowledge, ShortcutsBench is the first large-scale real API-based agent benchmark considering APIs, queries, and action sequences. ShortcutsBench provides rich and real APIs, queries with various difficulties and task types, high-quality human-annotated action sequences, and queries from real user demands. Moreover, it also provides precise values for parameter filling, including primitive data types, enum types, and the use of output from previous actions for parameter values, as well as evaluations of the agent’s awareness in requesting necessary input from the system or user. Furthermore, the scale of APIs, queries, and the corresponding action sequences is comparable or even better to benchmarks / datasets created by LLM or modified by existing datasets. The overall comparison between ShortcutsBench and existing benchmarks / datasets is listed in Table 2.
To demonstrate ShortcutsBench’s advantages, we do extensive evaluations of API-based agents from $10$ leading open-source and close-source LLMs, covering varying intelligence levels. To our best known, this is the most comprehensive evaluation considering the API selection, parameter value filling, and recognition of the need for input from the system or the user, covering all key processes of API-based agent. The evaluation results highlight great limitations of existing API-based agents.
In summary, this paper makes the following key contributions:
  • We identified problems of the existing benchmarks / datasets, specifically that they struggle to accommodate the advanced reasoning capabilities of existing more intelligent LLMs, and have conducted experiments to validate the problem.
  • We innovatively extracted data from Shortcuts, to build a high-quality benchmark for API-based agents. To our best knowledge, ShortcutsBench is the most realistic, rich, comprehensive, and large-scale benchmark for API-based agents. We hope this approach to dataset construction will inspire more researchers.
@@
  • We made efforts to evaluate $10$ advanced LLM-based agents with varying intelligence levels on the whole process required to complete user queries, including API selection, parameter filling, and their awareness of requesting necessary input from the system or user.
@@
  • We obtained massive interesting conclusions such as (1) Open-source LLM agents now match closed-source ones on simpler tasks but still lag behind on complex ones; (2) Extracting necessary parameters from queries is the most challenging task in parameter filling; (3) There is a substantial lack of awareness in agents when it comes to requesting the necessary input;
  • We have fully open-sourced all the datasets, code, experimental logs, and results, and provided detailed documents. We hope our research opens new directions for the real-world deployment of existing LLM-based agents.
@@
Table 2: ShortcutsBench has a great advantage in the ①realness and richness, ②the complexity of APIs, queries, and corresponding action sequences, ③the validity of action sequences, ④detailed parameter value filling, ⑤the awareness for asking necessary input, and ⑥the overall scale.
@@
# Apps 88 N/A 3451 3 N/A N/A 8 N/A N/A
# APIs 1414 390 16464 1645 53 400 232 13 464
# Queries 7627 21112 12657 17002 3938 274 2726 1530 18770
# Avg APIs 9.62 1.02 2.3 1.0 1.0 2.1 5.4 3.5* 2.65
# Avg Actions 21.62 1.02 4.0 1.0 1.0 2.2 5.9 3.9* 2.67
    denotes estimation.
@@ 2 Related Work
API-based agents treat APIs as tools. They accept queries, generate action sequences based on queries and provided APIs, and generate next action depends on the history actions (Wang et al., 2024b; Yao et al., 2023). Related work about API-based agents can generally be categorized into $3$ types: (1) Task-specific enhancement focuses on improving the agent’s ability like using the model (Shen et al., 2024; Zhong et al., 2024). (2) Data-driven workflows emphasize the importance of data by researching how to construct action sequences, enabling generated data to fine-tune the model (Qin et al., 2024; Patil et al., 2024). (3) Agent evaluation studies the assessment of agents (Huang et al., 2024b; Li et al., 2023).
@@
Code-based agents use code generated for interaction with the external environment. They accept queries, generate scripts in programming languages such as Python (OpenAI, 2024b; Wang et al., 2024c), JavaScript (Wang et al., 2024a; Zheng et al., 2023), or Shell (OpenInterpreter, 2024; Sladić et al., 2024), and then input the code into interpreters. The execution results are then returned to the agent, which is used to help determine the next code generation. Currently, these approaches primarily focus on enhancing agent performance in specific tasks by incorporating additional knowledge (Wang et al., 2024a; Wu et al., 2023), increasing feedback (OpenInterpreter, 2024; Huang et al., 2024a), and decomposing tasks (Huang et al., 2023; Prasad et al., 2024). In addition to work on optimization methods, numerous efforts have emerged to evaluate code-based agents (Trivedi et al., 2024; Liu et al., 2024)
@@
DAPs (Abdou et al., 2021) refer to software tools or services designed to optimize workflows through automation. DAPs leverage technologies such as robotic process automation (RPA) (Chakraborti et al., 2020) and low-code / no-code development tools to achieve the goals. DAPs like Zapier (Zapier, 2024), Make (Make, 2024), and IFTTT (Rahmati et al., 2017) offer extensive APIs that enable users to create automated workflows. Similarly, DAPs such as Microsoft Power Automate (Microsoft, 2024) and Tasker (Dias, 2024) are primarily used to build workflows on Azure and Android, respectively. Recently, with the rise of LLM-based agents, platforms like Coze (Coze, 2024) and Dify (Dify, 2024) have emerged as “agent construction platforms”. Functionality like “workflow” in these platforms can also help manually build workflows, but they have been specifically optimized for integration with LLMs.
Shortcuts app (formerly Workflow(Apple, 2024) is an app developed by Apple for building workflows through a graphical interface, available on Apple’s operating systems (iOS / iPadOS and macOS). Shortcuts app can be seen as the DAP of Apple. It allows users to create workflows (known as shortcuts (Apple, 2024c)) that execute specific tasks on their devices and share them online via iCloud link (Apple, 2024b). Users can also download curated shortcuts from the Gallery of the Shortcuts app. However, the shortcuts available in the Gallery are very limited, with only a few dozen options. To access more shortcuts, users must either collect them from third-party sharing sites like Shortcuts Gallery (Gallery, 2024) or create their own. Shortcuts can be triggered through the Shortcuts app, widgets, the share sheet, old Siri (Apple, 2024a), new Siri of Apple Intelligence (Apple, 2024b), and they can also be automated to run upon specific events.
Shortcuts are composed of multiple API calls (actions). An agent can use the shortcut as a whole API or utilize the individual APIs involved in the shortcut. This paper treats APIs within the shortcuts as APIs available to the agent, aiming for the agent to automatically construct workflows of API calls.
@@ 3 Dataset
In this Section, we first introduce the acquisition of the dataset (Section 3.1). Then, we outline the ShortcutsBench’s construction process (Section 3.2). Finally, we outline the setup for evaluation tasks to evaluate the agent’s ability to handle tasks of varying difficulty, including the ability to select suitable APIs (Section 3.3.1), the ability to do parameter filling (Section 3.3.2), and the awareness in requesting additional input from the system or user (Section 3.3.3).
@@ 3.1 Dataset Acquisition
Figure 1: (a) illustrates the data acquisition process. (b) shows the dataset acquisition of existing work. APIs in existing work are collected from API hubs, created by hand, or modified from existing datasets. The queries and action sequences are constructed using templates or semi / fully automated methods.
Figure 1 shows the data acquisition process. We first use search engines to identify popular public shortcut-sharing sites ①. We totally find $14$ sites (Table LABEL:tab:collected_14_sites). Then we crawled these sites to obtain fields such as “shortcut name”, “function description”, “shortcut type”, and “iCloud link” ②. Then we downloaded the shortcut source file by “iCloud link” and then perform deduplicating based on iCloud link itself and the actual shortcut content (i.e., the action sequences) ③. Subsequently, we extracted “app name” using the field WFWorkflowActionIdentifier in the shortcut source file, and then downloaded associated apps ④. These apps may come from various sources. (1) apps from the macOS or iOS App Store, (2) apps like Keynote from path /Applications/ and /System/Application/ on macOS, (3) third-party apps from the the official website of the app. During the downloading, we also excluded some legacy apps and paid apps.
Then we managed to extract APIs from the downloaded apps ⑤. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Apple-Inc., 2024b) framework and \${filename}.intentdefinition from SiriKit (Apple-Inc., 2024c) framework. We extracted all APIs involved in the apps. During the extraction, we perform deduplication of APIs based on manually crafted rules as an app may have multiple duplicate API definition files with the same API definition. This process also involves significant manual filtering. Additionally, for app Shortcuts, which are deeply integrated with Apple’s operating system, we need to obtain their API definition files WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS, instead of extracting it from the app itself. Subsequently, we further filtered the shortcuts ⑥ based on criteria such as whether the associated apps is paid app, whether the apps were outdated, and whether the APIs were deprecated. Additionally, we imported all shortcuts into the macOS Shortcuts app to ensure they were functional. Finally, as shown in Table 2, we get $88$ apps from various categories such as “Health & Fitness” and “Developer Tools”. We finally get $1414$ APIs involved in $7627$ shortcuts.
As the acquisition process involves specific knowledge about shortcut-sharing sites and Apple’s operating system, detailed explanations are omitted here due to space constraints. For more details about the whole acquisition process, please refer to Appendix A.1.